OpenStudy (anonymous):

What is the distance between P(–4, 3) and Q(6, 1)? Round to the nearest tenth?

OpenStudy (saifoo.khan):

Do u know the distance formula?

OpenStudy (anonymous):

?no do you?

OpenStudy (saifoo.khan):

Yes i do. wanna know?

OpenStudy (anonymous):

?no do you?

OpenStudy (anonymous):

yes please

OpenStudy (saifoo.khan):

http://0.tqn.com/d/math/1/0/A/1/distance.gif

OpenStudy (anonymous):

?no do you?

OpenStudy (anonymous):

o.o sorry it resent the last thing

OpenStudy (anonymous):

\[PQ=\sqrt{(3-1)^2+(6+4)^2}\]\[=\sqrt{4+100}=\sqrt{104}\]\[=2\sqrt{26} \approx10.2\]

OpenStudy (anonymous):

right ten units, down 2 units so by pythagoras your have a right triangle with hypotenuse \[\sqrt{10^2+2^2}\]

OpenStudy (anonymous):

thanks!!!!

OpenStudy (anonymous):

page 52